The function P=11+44lnx represents the percentage of spam email where x is the number of years after 2002. What is the approximate percentage of spam mail in the year 2006 to the nearest tenth of a percent.Calculate to six decimals when necessary.Please explain how you obtained this answer.
Answer by ankor@dixie-net.com(22740)

he function P=11+44lnx represents the percentage of spam email where x is the number of years after 2002. What is the approximate percentage of spam mail in the year 2006 to the nearest tenth of a percent.Calculate to six decimals when necessary.Please explain how you obtained this answer.
:
x years after 2002, 2006: x = 4
:
P = 11 + 44*ln(4)
:
Find the ln of 4
P = 11 + 44(1.286294)
:
P = 11 + 60.996952
:
P = 71.996952 ~ 72.0% is spam in 2006